Obtain PDE z = x + ax2 y2 where a and b are arbitrary constants and also

analyze the partial differential equation. 


1
Expert's answer
2021-02-24T07:37:51-0500

Given equation is-

"z=x+ax^2y^2"


Differentiate w.r.t. to x partially


"\dfrac{dz}{dx}=1+2axy^2"


"(\dfrac{dz}{dx}-1)=2axy^2"


"a=\dfrac{1}{2xy^2}(\dfrac{dz}{dx}-1)"


Putting the value of a in given equation-

"z=x+\dfrac{1}{2xy^2}(p-1)x^2y^2" {"\dfrac{dz}{dx}=p" }


="x+\dfrac{1}{2}(p-1)x"


"=x+\dfrac{px}{2}-\dfrac{x}{2}"


z= "\dfrac{x(1+p)}{2}"